Put in Y form x+4=y/2+3
Answer by algebrahouse.com(1659)

x + 4 = (y/2) + 3
x + 1 = y/2 {subtracted 3 from both sides}
y = 2(x + 1) {multiplied both sides by 2}
y = 2x + 2 {used distributive property}

For more help from me, visit: www.algebrahouse.com